If a thirty-year standard coupon bond has a coupon rate of 6.8%, and the YTM is 6.4% then what would be the quarterly coupons?
Answer #1

Suppose the face value=$1000
Coupon rate=6.8%
Quarterly coupon rate=6.8%/4=0.017

Quarterly coupon payment=(Quarterly coupon rate)*(Face value)=0.017*1000=17
So, the quarterly coupons=$17